OpenStudy (anonymous):

Solve for x: square root of (x+4)=x-2. I'm trying to do this problem and I think I did something wrong but can't tell what it is. Right now I have: x^2-x=0

OpenStudy (anonymous):

Is it √ ( x +4) = x -2 ?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Have you ( x-2)² = ....

OpenStudy (anonymous):

Yes. And, I have \[x+4=x^2+4\]

OpenStudy (anonymous):

( x-2)² = x² - 4x + 4

OpenStudy (anonymous):

Since ( x-2)² = ( x -2) ( x -2)

OpenStudy (anonymous):

Oooooooooooh. Right. I always forget to FOIL...

OpenStudy (anonymous):

Can you get this result? ( x-2)² = x² - 4x + 4

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

After working it out, I now have\[x^2-5x\]

OpenStudy (anonymous):

Does x=0?

OpenStudy (anonymous):

x ( x - 5) = 0

OpenStudy (anonymous):

Can you solve for x?

OpenStudy (anonymous):

x=0, 5?

OpenStudy (anonymous):

Is that my final answer or can I only have 1?

OpenStudy (anonymous):

Alright. Thank you! :)

OpenStudy (anonymous):

@alienshe I forgot to check x-2 ≥ 0 => x ≥ 2 Thus only one solution x = 5 accepted